OpenStudy (anonymous):

Simplify the following expression 5 2/3 divided by 11 2/5 a.85/171 b.1/15 c16 68/77 d.113/130

jabez177 (jabez177):

First we need to convert the Whole fractions into improper fractions.

jabez177 (jabez177):

5 x 3 + 2 = ?

OpenStudy (anonymous):

17/3 divided by 57/5

jabez177 (jabez177):

Good. And when we are dividing fractions, it's the same as multiplying the reciprocal. So \[\frac{ 17 }{ 3 }\times \frac{ 5 }{ 57 }\]

jabez177 (jabez177):

17 x 5 is what?

OpenStudy (anonymous):

85?

jabez177 (jabez177):

Good, now what is 3 x 57?

OpenStudy (anonymous):

225

jabez177 (jabez177):

Nope, it's 171 :)

jabez177 (jabez177):

So it's \[\frac{ 85 }{ 171 }\]

jabez177 (jabez177):

Can it be simplified further?

OpenStudy (anonymous):

thank you very much

jabez177 (jabez177):

You're welcome! Click on "Best Response" once you are satisfied with your answer. :) Welcome to OpenStudy! :D